May 30, 2014

Observer: Zheng Cai, Yuguang Chen, Ian McGreer

Comments: Clear all night, seeing 0.7- 1.0''

Time (MST):

8:20pm Dome open.

8:20- 8:50pm Collimation problem, being solved

8:50- 9:00pm Aquisition on Cai's mask

9:00- 1:20am Finish 8x30min on Cai's mask

1:25- 2:05am run McGreer 's MODS imaging script

2:05- 2:20am Actuator force on mirror exceed limit, go to zenith, re-collimation.

2:25- 2:45am Aqusition on Smith's target SN2008S,

2:45- 3:07am finish 1x20min (last night, seeing getting bigger on the last 20-min spectrum. Half of the last spectrum was also taken during the twilight. Since this target is high priority, just add 20min on this target).

3:07- 3:28am aqusition and try to match the target with the finding chart. It seems this target gets faint by a lot.

3:28- 4:00am finish 2x15min exposure, faint continua.

Edited by Zheng Cai

-- %USERSIG{EiichiEgami - 2014-05-26}%

Comments


 
Topic revision: r2 - 31 May 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback